OpenStudy (anonymous):

simplify the trigonometric expression sin^2 theta/1-cos theta

OpenStudy (anonymous):

Multiply the top and bottom by 1+costheta

OpenStudy (anonymous):

can you show me how

OpenStudy (anonymous):

\[\frac{ \sin^2\theta }{1-\cos \theta }*\frac{ 1+\cos \theta }{ 1+\cos \theta }\]

jimthompson5910 (jim_thompson5910):

sin^2 = 1-cos^2 then use the difference of squares rule to factor

OpenStudy (anonymous):

so sin^2 theta/ -1 -cos theta?